Russell Preston delivers parts for several local auto parts stores. He charges clients $1.20 per mile driven. Russell has determined that if he drives 2,700 miles in a month, his average operating cost is $0.80 per mile. If he drives 3,700 miles in a month, his average operating cost is $0.70 per mile. Russell has used the high-low method to determine that his monthly cost equation is total cost = $1,540 + $0.43 per mile. Required:

1. Determine how many miles Russell needs to drive to break even.

2. Assume Russell drove 2,500 miles last month. Without making any additional calculations, determine whether he earned a profit or a loss last month.

3. Determine how many miles Russell must drive to earn $1,925 in profit.

4-a. Prepare a contribution margin income statement assuming Russell drove 2,500 miles last month.

4-b. Use the information provided in Req 4a to calculate Russell’s degree of operating leverage.
Answer #1
1
Fixed costs 1540
Divide by Unit Contribution margin 0.77 =1.20-0.43
Break-even miles 2000
2
Profit, as miles is above 2000
3
Fixed costs 1540
Add: Profit 1925
Required Contribution margin 3465
Divide by Unit Contribution margin 0.77 =1.20-0.43
Target miles 4500
4a
Contribution margin income statement
Sales revenue 3000 =2500*1.20
Variable costs 1075 =2500*0.43
Contribution margin 1925
Fixed costs 1540
Income from Operations 385
b
Contribution margin 1925
Divide by Income from Operations 385
Degree of Operating leverage 5.00